Given the function

𝑔(π‘₯) = π‘₯ 𝐽1 (π‘₯) βˆ’ 𝐴 𝐽0(π‘₯) with 𝐴 = π‘π‘œπ‘›π‘ π‘‘π‘Žπ‘›π‘‘ β‰₯ 0,


determine 𝑔′, 𝑔 β€²β€² and 𝑔′′′ . Reduce all expressions to functions of 𝐽0(π‘₯), 𝐽1(π‘₯) and 𝐴 only.



1
Expert's answer
2021-05-11T09:18:57-0400

"g(x)=xJ_1(x)-AJ_0(x)" ....1)


Now differentiating the above equation we get-


"g^{'}(x)=J_1(x)-xJ_1^{'}(x)-AJ_0^{'}(x)"



Now by the recurrence relation ,we know that


"J_n^{'}(x)=\dfrac{J_n-1(x)-J_n+1(x)}{2}" ......2)


now putting n=1 and n=0 we get the values of "J_1^{'}(x)\ and \ J_0^{'}(x)."


now putting the values we get g"^{'}(x)=J_1(x)-x \dfrac{J_0{(x)}-J_2(x)}{2}-A \dfrac{J_{-1}(x)-J_1(x)}{2}" which is our required equation.



"g^{''}(x)=J_1^{'}(x)-J_1^{'}(x)-xJ_1^{''}(x)-AJ_o^{''}(x)"


Now differentiating recurrence relation .....2)


"J_n^{"}(x)= \dfrac{J_n-1^{'}(x)-J_n+1{'}(x)}{2}" .....3)

Now in recurrence relation .....2), if we replace n"\to" n-1,we get -"J_{n-1}^{'}(x)=\dfrac{J_{n-2}(x)-J_n(x)}{2}"

now replacing n"\to" n+1, we get- "J_n+1{'}(x)=\dfrac{Jn(x)-J_{n+2}}{2}"

now puttting both these equation in .....3), we get-"\ Jn{''}(x)=" "\dfrac{J_{n-2}(x)-2J_n(x)-J_{n+2}(x)}{2}"


now putting n=1,0 we get -


"g^{''}(x)=-x\dfrac {J_{-1}(x)-2J_1{(x)}}{2}"